OpenStudy (anonymous):

Can someone check my answer to see if it is correct? What is 'f' and 'a' of the function. lim as x approaches pi/4 (tanx-1)/(x-(pi/4)) a= pi/4 f=tan x

OpenStudy (anonymous):

yes, you are being asked for the derivative of \(\tan(x)\) at \(x=\frac{\pi}{4}\) your answer is correct

OpenStudy (anonymous):

\[\tan'(\frac{\pi}{4})=\lim_{x\to \frac{\pi}{4}}\frac{\tan(x)-\tan(\frac{\pi}{4})}{x-\frac{\pi}{4}}\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yw